### v1

@@ 1 Introduction to the Universe
For millennia humans have gazed upon the night sky and pondered what our place is in the Universe and what laws govern it? The unaided human eye can only see about 6000 stars but for many years their distances, lives and death were a mystery. It wasn’t until astronomers applied mathematics to observational data that our understanding of the Universe and the night sky started to progress rapidly. Johannes Kepler took Tycho Brahe’s precision measurements of the positions of the planets to understand their trajectories, and developed his famous three laws of planetary motion. A further revolution in our understanding of the Universe occurred when Galileo made the first recorded observations of the night sky with a telescope in the early 1600s. His observations revealed that Jupiter was orbited by four major moons, and this displaced the prevailing view that Earth was the centre of the Universe. Four centuries of observations have followed, and the astronomical and theoretical breakthroughs are astonishing. The heavens have changed from a static background with pinpricks of light and wandering planets to a dynamic and violent Universe where phenomenal instruments are revealing the innermost workings of complex systems involving stars, planets, black holes, white dwarfs and neutron stars.
@@ 2 The transient radio sky
Although optical astronomy has existed since humans first gazed at the heavens, radio astronomy is less than 100 years old. The first radio telescopes were not sensitive to short timescale phenomena; they focused on making maps of the radio sky that were unchanging and static. With the discovery of quasars in 1964, however, that changed. Because the angular sized of quasars are so small, quasars are observed to “scintillate”, or twinkle as their radio waves travel through interstellar gas. This effect, due to diffraction of the waves through the interstellar medium, causes their radio brightness to fluctuate with time.
@@
Soon Bell Burnell and Hewish found more of these objects, in different parts of the sky, confirming that they were a new class of astronomical sources. Following the suggestion of a journalist, these “pulsating radio sources” were henceforth referred to as pulsars. Not long after, Thomas Gold and Franco Pacini independently suggested that these pulses were neutron stars born in the supernova explosions of massive stars. Soon thereafter, when the 33-ms Crab pulsar was discovered at the centre of the Crab supernova remnant in 1968 [63] and the 89-ms pulsar was found at the centre of the Vela supernova remnant [42], this theory was confirmed.
@@
Following a number of attempts to find extragalactic radio pulses at radio wavelengths [22, 35, 32, 14, 55, 25], Linscott and Erkes [44] announced the discovery of radio pulses detected from the direction of the giant elliptical galaxy M87. However, in spite of a number of follow-up searches [65, 29, 50, 64], no further radio pulses were ever seen from M87. While it was never conclusively determined whether the events seen by Linscott and Erkes were from highly variable sources in M87 or some form of interference, we note that the galaxy is strongly suspected to contain detectable sources [41]. Large-area searches for extragalactic radio pulses in the intervening years continued to yield null results [1, 7, 36]. Meanwhile, Palmer [54] had emphasized the importance of dispersion in the detection of cosmological radio transients and searches for ‘prompt’ radio emission from gamma-ray bursts were being carried out [see, for example, 28, 4]
@@ 5 The repeating FRB — a “minor point of interest”
The repetition of this FRB made it possible to observe its position with an interferometer to localise it much more precisely than possible with a single-dish telescope like Arecibo. Observations with the VLA by Chatterjee et al [16] revealed that the burst was associated with a galaxy. This was very important for several reasons. First, these bursts were definitely astrophysical and coming from stellar populations. Second, as shown in the figure below, the galaxy was detected optically by Tendulkar et al [66], providing a measurement of the redshift, $z=0.19$. Hence, for the first time the DM of a burst could be used to constrain the electron density of the intergalactic medium. Third, the kind of galaxy it came from provided important clues to its origin. In this case, FRB 121102 originated in a dwarf galaxy with a high star formation rate, indicating that the burst came from stellar populations such as neutron stars. It was also found that the bursts from FRB 121102 had one of the highest rotation measures ever observed [53] which means it came from a very magnetic, dense environment.
@@ 6 Recent breakthroughs and the future
Regardless of all the mysteries that are still in play, one thing is very clear: as we look at the current state of the transient sky in terms of the “luminosity — pulse duration” diagram shown below, we see clear groupings of sources, but also lots of gaps in the parameter space.
@@
These gaps indicate opportunities and challenges to the FRB community. We are not, for example, currently able to resolve pulses well enough to rule out a class of extremely short duration FRBs (dubbed ultrafast FRBs by some). Similarly, we are currently less sensitive to long-duration pulses which may turn out to be extremely numerous, particularly in the Milky Way where examples of long period (minutes or more) transients are now being found [33, 34]. It is important to recognize these current shortcomings and opportunities as we plan future experiments. There is still much to explore.
@@ Acknowledgements
We are grateful to a number of individuals for their roles in this work and organisations for supporting it over the last 17 years. Special thanks to Froney Crawford and Ash Narkevic for their contributions to the discovery. To the Shaw Foundation, and to those that nominate awardees and serve on our committees, thank you for highlighting the significance of the discovery of FRBs. Matthew has received significant support from Swinburne University of Technology and the Australian Research Council. Maura and Duncan are grateful for the sustained support from WVU, the National Science Foundation, the Research Corporation for Science Advancement and the West Virginia NASA Space Grant Consortium as well as the West Virginia Higher Education Policy Commission. We are grateful to our colleagues at the Australia Telescope National Facility, and the National Radio Astronomy Observatory for their support. Duncan would like to thank Enrico Massaro and Frank Drake for alerting him to some of the key references and developments concerning searches for extragalactic radio pulses prior to the discovery of FRBs. We are very grateful to Fabio Santos and Elias Brinks at Springer for their patience as we put this manuscript together, and to the reviewer, Vicky Kaspi, for very useful feedback to an earlier draft of the manuscript. Finally, huge thanks to all of our students and colleagues who have been a pleasure to work with and a constant source of inspiration.

### v3

@@ 1 Introduction to the Universe
For millennia humans have gazed upon the night sky and pondered what our place is in the Universe and what laws govern it? The unaided human eye can only see about 6000 stars but for many years their distances, lives and death were a mystery. It wasn’t until astronomers applied mathematics to observational data that our understanding of the Universe and the night sky started to progress rapidly. Following the Copernican Revolution which inspired Tycho Brahe’s precision measurements of the positions of the planets, Johannes Kepler used Brahe’s work to understand their trajectories, and developed his famous three laws of planetary motion. A further revolution in our understanding of the Universe occurred when Galileo made the first recorded observations of the night sky with a telescope in the early 1600s. His observations revealed that Jupiter was orbited by four major moons, and this displaced the prevailing view that Earth was the centre of the Universe. Four centuries of observations have followed, and the astronomical and theoretical breakthroughs are astonishing. The heavens have changed from a static background with pinpricks of light and wandering planets to a dynamic and violent Universe where phenomenal instruments are revealing the innermost workings of complex systems involving stars, planets, black holes, white dwarfs and neutron stars.
@@ 2 The transient radio sky
Although optical astronomy has existed since humans first gazed at the heavens, radio astronomy is less than 100 years old. The first radio telescopes were not sensitive to short timescale phenomena; they focused on making maps of the radio sky that were unchanging and static. With the discovery of quasars in 1964, however, that changed. Because the angular sizes of quasars are so small, quasars are observed to “scintillate”, or twinkle as their radio waves travel through the ionized interplanetary medium.
@@
Soon Bell Burnell and Hewish found more of these objects, in different parts of the sky, confirming that they were a new class of astronomical sources. Following the suggestion of a journalist, these “pulsating radio sources” were henceforth referred to as pulsars. Not long after, Thomas Gold and Franco Pacini both suggested that these pulses were neutron stars born in the supernova explosions of massive stars. Soon thereafter, when the 33-ms Crab pulsar was discovered at the centre of the Crab supernova remnant in 1968 [65] and the 89-ms pulsar was found at the centre of the Vela supernova remnant [44], this theory was confirmed.
@@
Following a number of attempts to find extragalactic radio pulses at radio wavelengths [24, 37, 34, 16, 57, 27], Linscott and Erkes [46] announced the discovery of radio pulses detected from the direction of the giant elliptical galaxy M87. However, in spite of a number of follow-up searches [67, 31, 52, 66], no further radio pulses were ever seen from M87. While it was never conclusively determined whether the events seen by Linscott and Erkes were from highly variable sources in M87 or some form of interference, we note that the galaxy is strongly suspected to contain detectable sources [43]. Large-area searches for extragalactic radio pulses in the intervening years continued to yield null results [1, 9, 38]. Meanwhile, Palmer [56] had emphasized the importance of dispersion in the detection of cosmological radio transients and searches for ‘prompt’ radio emission from gamma-ray bursts were being carried out [see, for example, 30, 4]
@@ 5 The repeating FRB — a “minor point of interest”
The repetition of this FRB made it possible to observe its position with an interferometer to localise it much more precisely than possible with a single-dish telescope like Arecibo. Observations with the VLA by Chatterjee et al [18] revealed that the burst was associated with a galaxy. This was very important for several reasons. First, these bursts were definitely astrophysical and coming from stellar populations. Second, as shown in Fig. 7, the galaxy was detected optically by Tendulkar et al [68], providing a measurement of the redshift, $z=0.19$. Hence, for the first time the DM of a burst could be used to constrain the electron density of the intergalactic medium. Third, the kind of galaxy it came from provided important clues to its origin. In this case, FRB 121102 originated in a dwarf galaxy with a high star formation rate, indicating that the burst came from stellar populations such as neutron stars. It was also found that the bursts from FRB 121102 had one of the highest rotation measures ever observed [55] which means it came from a very magnetic, dense environment.
@@ 6 Recent breakthroughs and the future
Regardless of all the mysteries that are still in play, one thing is very clear: as we look at the current state of the transient sky in terms of the “luminosity — pulse duration” diagram shown in Fig. 9, we see clear groupings of sources, but also lots of gaps in the parameter space.
@@
These gaps indicate opportunities and challenges to the FRB community. We are not, for example, currently able to resolve pulses well enough to rule out a class of extremely short duration FRBs (dubbed ultrafast FRBs by some). Following predictions about a substantial population of ultra-long period Galactic magnetars [7], and the fact that most searches to date are less sensitive to long-duration pulses, examples of long period (minutes or more) transients in the Galaxy are now being found [35, 36]. It is important to recognize these current shortcomings and opportunities as we plan future experiments. Further discussion of this emerging population can be found in Beniamini et al [8].
@@ Acknowledgements
We are grateful to a number of individuals for their roles in this work and organisations for supporting it over the last 17 years. Special thanks to Froney Crawford and Ash Narkevic for their contributions to the discovery. To the Shaw Foundation, and to those that nominate awardees and serve on our committees, thank you for highlighting the significance of the discovery of FRBs. Matthew has received significant support from Swinburne University of Technology and the Australian Research Council. Maura and Duncan are grateful for the sustained support from WVU, the National Science Foundation, the Research Corporation for Science Advancement and the West Virginia NASA Space Grant Consortium as well as the West Virginia Higher Education Policy Commission. We are grateful to our colleagues at the Australia Telescope National Facility, and the National Radio Astronomy Observatory for their support. Duncan would like to thank Enrico Massaro and Frank Drake for alerting him to some of the key references and developments concerning searches for extragalactic radio pulses prior to the discovery of FRBs, Paz Beniamini for highlighting the importance of long-period pulsars in the Galaxy and Ron Ekers for helpful comments. We are very grateful to Fabio Santos and Elias Brinks at Springer for their patience as we put this manuscript together, and to the reviewer, Vicky Kaspi, for very useful feedback to an earlier draft of the manuscript. Finally, huge thanks to all of our students and colleagues who have been a pleasure to work with and a constant source of inspiration.